OpenStudy (anonymous):

The length of a rectangle is 2 meters more than its width.If the area is 63 meters,what are the length and width?

OpenStudy (amistre64):

Area = LW; L=W+2 63 = (W+2)(W) 63 = W^2 + 2W W^2 +W - 63 = 0 so far so good?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

the fastest way to solve a quadratic is with the quadratic formula, do you know the quadratic formula?

OpenStudy (anonymous):

not really

OpenStudy (amistre64):

Do you know how to "complete the square"? which is the long proof for the quad formula?

OpenStudy (anonymous):

no

OpenStudy (amistre64):

which one would you like to learn?

OpenStudy (anonymous):

the fastest way

OpenStudy (amistre64):

lol.... good answer

OpenStudy (amistre64):

the quad formula is: -(middle) +- sqrt(middle^2 - (4)(first)(last)) -------------------------------------- 2(first) Lets plug it in :)

OpenStudy (amistre64):

(1)W^2 +(1)W - 63 -1 +- sqrt(1^2 -4(1)(-63)) ------------------------ 2(1)

OpenStudy (amistre64):

-1 +- sqrt(1+252) ---------------- 2

OpenStudy (amistre64):

-1 +- sqrt(253) --- -------- 2 2 now we use a calculator :)

OpenStudy (amistre64):

W = about 7.45

OpenStudy (amistre64):

we actually get 2 answers, but we can throw out the negative because area is not a negative measurement

OpenStudy (amistre64):

L = W+2 so about 9.45 if I did it right :)

OpenStudy (amistre64):

becuase I chopped off alot of decimals, we get a skewed result. but its close

OpenStudy (anonymous):

so we divided the number above and this is the results

OpenStudy (amistre64):

as close as I can get :) barring I havent made any mistakes

OpenStudy (anonymous):

so what is the square root of 253

OpenStudy (anonymous):

i got 7.95

OpenStudy (amistre64):

sqrt(253)=15.905973720586866353745477830541

OpenStudy (anonymous):

yes and divided by 2 is 7.95529865

OpenStudy (amistre64):

-.5

OpenStudy (anonymous):

7.95 + 2 meters is 9.95 correct

OpenStudy (amistre64):

still need to take away .5 from 7.955.......

OpenStudy (anonymous):

where did -5 come from

OpenStudy (amistre64):

-1 +- sqrt(253) --- -------- 2 2 the first part here is -1/2

OpenStudy (anonymous):

ohhhhhh.....ok.......i see

OpenStudy (anonymous):

so what does the final answer look like

OpenStudy (amistre64):

itll probably include a radical sign for an eexat nswer; but the best I can do with an estimate is: W=about 7.45 and L=about 9.45

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

can you help me with another one

OpenStudy (amistre64):

i dunno, my computer is slowing down something feirce. try it in another question post; and ill see what i can do

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

I know...iknow..... i see where I messed up. A=63; L=9 and W=7 i lost a "2" in the equation somewhere lol... w^2 +[2]w -63 = 0 <---THAT ONE!! -2/2 + sqrt(4 - (4)(1)(-63))/2 -1 + sqrt([252] +4)/2 <-AND I forgot how to multiply -1 + sqrt(256)/2 -1 + 16/2 -1 + 8 W = 7 L= W+2 = 7+2 = 9 Area = 9(7) = 63 Are ya happy now :)